# Dual Consolidation for Pre-Trained Model-Based Domain-Incremental Learning

### v1

@@ Abstract
Domain-Incremental Learning (DIL) involves the progressive adaptation of a model to new concepts across different domains. While recent advances in pre-trained models provide a solid foundation for DIL, learning new concepts often results in the catastrophic forgetting of pre-trained knowledge. Specifically, sequential model updates can overwrite both the representation and the classifier with knowledge from the latest domain. Thus, it is crucial to develop a representation and corresponding classifier that accommodate all seen domains throughout the learning process. To this end, we propose DUal ConsolidaTion (Duct) to unify and consolidate historical knowledge at both the representation and classifier levels. By merging the backbone of different stages, we create a representation space suitable for multiple domains incrementally. The merged representation serves as a balanced intermediary that captures task-specific features from all seen domains. Additionally, to address the mismatch between consolidated embeddings and the classifier, we introduce an extra classifier consolidation process. Leveraging class-wise semantic information, we estimate the classifier weights of old domains within the latest embedding space. By merging historical and estimated classifiers, we align them with the consolidated embedding space, facilitating incremental classification. Extensive experimental results on four benchmark datasets demonstrate Duct’s state-of-the-art performance.
@@ 1 Introduction
Recent years have seen the rise of deep learning, demonstrating its strong potential in real-world applications He et al. (2016); Deng et al. (2009); Floridi & Chiriatti (2020). However, in a dynamic and ever-changing world, data often evolves in a stream format Aggarwal (2018); Ye et al. (2024), necessitating continuous updates with data from new domains Gama et al. (2014). For instance, autonomous vehicles must handle driving tasks across different seasons and weather conditions Bojarski et al. (2016), and face recognition systems are supposed to recognize users despite the varying lighting conditions Zhao et al. (2003). Correspondingly, Domain-Incremental Learning (DIL) van de Ven et al. (2022) addresses this challenge by absorbing new knowledge from new data distributions while preserving existing knowledge. However, sequentially updating the model with new domains often biases the embedding Yu et al. (2020) and classifier Zhao et al. (2020); Wu et al. (2019) modules toward the latest domain, leading to catastrophic forgetting Mermillod et al. (2013); De Lange et al. (2021) of previous knowledge. Therefore, developing methods for learning new domains without forgetting existing knowledge becomes a critical challenge for the machine learning community.
To combat catastrophic forgetting, recent works leverage strong pre-trained models (PTMs)Han et al. (2021) as the initialization for DILWang et al. (2022d; c; b). PTMs’ robust representation abilities provide a powerful starting point, showing promising results in DIL benchmarks. These approaches freeze the pre-trained backbone to preserve existing knowledge and append lightweight modules Jia et al. (2022); Chen et al. (2022); Lian et al. (2022) to capture new patterns. However, since incoming domains will also overwrite the appended modules, the representations still suffer from forgetting, leading to corresponding bias in the classifier.
In DIL, there are two primary sources of forgetting: the overwriting of features and classifiers. Continually adjusting the features to capture the latest task biases the representations toward the newest domain. For example, suppose the previous task involves real-world photos while the subsequent task includes quick-draw images. The features will be overwhelmingly adjusted to highlight non-objective details, which are less helpful in distinguishing the previous domain, thus leading to forgetting. In light of this, an ideal feature space in DIL should fit all domains and equally highlight the domain-specific features for all seen tasks. However, as features become less generalizable due to overwriting, the classifier also becomes biased toward the latest domain. Since the classifier plays the pivotal role of mapping embeddings to classes, using the biased classifier can cause detrimental effects to decision-making. Consequently, it is imperative to consolidate existing knowledge both in the feature and classifier to resist forgetting.
@@
In this paper, we propose DUal ConsolidaTion (Duct) to address these challenges. To counter feature drift, we introduce a representation merging technique that continuously integrates historical backbones. This merged embedding consolidates task-specific information from all previous domains, providing informative features without forgetting. Consequently, we achieve non-forgettable features incrementally and resist feature-level forgetting. To map updated features to classes, we design a classifier consolidation process that merges historical classifier weights with calibrated weights. By extracting class-wise semantic relationships in the joint space, we develop a classifier transport mechanism that estimates classifiers of previous domain classes in the latest embedding space. Through consolidating features and classifiers in this coordinated way, Duct balances all seen domains and robustly resists forgetting, achieving competitive performance on various benchmarks.
@@ 2 Related Work
Continual Learning/Incremental Learning is a popular topic in the machine learning field, aiming to incorporate new knowledge within streaming data De Lange et al. (2021); Masana et al. (2023); Wang et al. (2024); Zhao et al. (2021); Zheng et al. (2024); Qi et al. (2024). It can be further classified into three subcategories, i.e., task-incremental learning (TIL) De Lange et al. (2021), class-incremental learning (CIL) Masana et al. (2023), and domain-incremental learning (DIL) van de Ven et al. (2022). Among them, TIL and CIL are faced with emerging new classes and are required to learn them without forgetting. By contrast, DIL mainly focuses on the scenario where label space is fixed while incoming data involves new domains Wang et al. (2022b); Shi & Wang (2023). In continual learning, typical solutions include using data replay Bang et al. (2021); Aljundi et al. (2019b); Isele & Cosgun (2018); Rolnick et al. (2019); Ratcliff (1990) to review former knowledge, using knowledge distillation Hinton et al. (2015); Rebuffi et al. (2017); Li & Hoiem (2017) or parameter regularization Kirkpatrick et al. (2017); Zenke et al. (2017); Chaudhry et al. (2018); Aljundi et al. (2018; 2019a) to maintain existing knowledge. Recent advances also resort to bias correction Hou et al. (2019); Zhao et al. (2020); Ahn et al. (2021); Castro et al. (2018); Wu et al. (2019); Belouadah & Popescu (2019) to rectify the model bias or expand the network structure as data evolves Rusu et al. (2016); Aljundi et al. (2017); Yan et al. (2021); Wang et al. (2022a); Zhou et al. (2023).
Domain-Incremental Learning with Pre-Trained Models: With the rapid development of pre-training techniques, PTMs provide a strong initialization for DIL models to boost the feature generalizability Smith et al. (2023); Wang et al. (2022c; d; b). Current PTM-based DIL methods mainly resort to visual prompt tuning Jia et al. (2022) to adjust the pre-trained features. With the pre-trained features frozen, it learns a prompt pool as external knowledge and selects instance-specific prompts to encode task information. L2P Wang et al. (2022d) utilizes the query-key matching mechanism to select instance-specific prompts, while DualPrompt Wang et al. (2022c) extends it by learning task-specific and instance-specific prompts jointly. By contrast, S-Prompts Wang et al. (2022b) learns domain-specific prompts for DIL and retrieves prompts via KNN search. To alleviate the prompt selection cost, CODA-Prompt Smith et al. (2023) modifies the prompt selection process into an attention-based weighted combination. Apart from prompt tuning, there are also some methods that directly build classifiers upon the pre-trained features Zhou et al. (2024a); McDonnell et al. (2023).
Model Merging: Recent advances in pre-training have made weight interpolation among different models a useful technique in model editing Ainsworth et al. (2023); Frankle et al. (2020); Ilharco et al. (2023); Matena & Raffel (2021); Wortsman et al. (2022a; b). Most work focuses on simple averaging of multiple models, aiming to enhance model in terms of its performance on the single task and robustness to out-of-distribution data. In this paper, we consider merging the pre-trained weights and task vector parameters to cultivate multitask representation for all seen domains.
@@ 3 From Old Domains to New Domains
In this section, we introduce the background information about domain-incremental learning and the baselines for applying pre-trained models to domain-incremental learning.
@@ 3.1 Domain-Incremental Learning
In domain-incremental learning, the model is faced with the data stream with a sequence of tasks $\left\{\mathcal{D}^{1},\mathcal{D}^{2},\cdots,\mathcal{D}^{B}\right\}$, where $\mathcal{D}^{b}=\left\{\mathcal{X}_{b},\mathcal{Y}_{b}\right\}$ is the $b$-th training task, i.e., $\mathcal{X}_{b}=\left\{{\bf x}_{i}\right\}_{i=1}^{n_{b}}$ and $\mathcal{Y}_{b}=\left\{y_{i}\right\}_{i=1}^{n_{b}}$. Each training instance ${\bf x}_{i}\in\mathbb{R}^{D}$ belongs to class $y_{i}\in Y$. This paper follows the exemplar-free setting Zhu et al. (2021); Wang et al. (2022d), i.e., during the $b$-th training stage, we can only access data in the current dataset $\mathcal{D}^{b}$. In DIL, the label space $Y$ is unchanged throughout the learning process, while the distribution of input instance keeps changing from domain to domain, i.e., $p(\mathcal{X}_{b})\neq p(\mathcal{X}_{b^{\prime}})$ for $b\neq b^{\prime}$. The target of DIL is to fit a model $f({\bf x})$ that can discriminate the classes among any seen domains:
@@
Following Wang et al. (2022c; d; b); Smith et al. (2023), we assume a pre-trained Vision Transformer (ViT) Dosovitskiy et al. (2020) is available as the initialization for $f({\bf x})$. For simplicity, we decouple the network structure into the embedding function $\phi(\cdot):\mathbb{R}^{D}\rightarrow\mathbb{R}^{d}$ (i.e., the final [CLS] token) and the linear classifier $W\in\mathbb{R}^{d\times|Y|}$. In this way, the output is denoted as $f({\bf x})=W^{\top}\phi({\bf x})$, and we utilize a cosine classifier in this paper. The classifier is further decoupled into $W=[{\bf w}_{1},{\bf w}_{2},\cdots,{\bf w}_{|{Y}|}]$, where ${\bf w}_{j}$ denotes the classifier weight of the $j$-th class. Although the label space is static during the learning process, data distribution of the same class across different domains can yield significant domain gaps. Hence, we follow existing works Wang et al. (2022c; d; b); Smith et al. (2023) to expand the classifier $W\in\mathbb{R}^{d\times b|Y|}$ as new tasks emerge, and the final prediction is made by $\left(\operatorname*{arg\,max}_{i}{\bf w}_{i}^{\top}\phi({\bf x})\right)\textbf{ mod }|Y|$.
@@ 3.2 Baselines in Domain-Incremental Learning
In domain-incremental learning, a naive solution facing the incoming datasets of new domains is to directly optimize the embedding and classifier:
@@
With pre-trained weights as initialization, Eq. 2 can quickly capture the discriminative features within the new domain. However, since the embedding $\phi(\cdot)$ is kept changing among different domains, it quickly loses the generalization ability to previous domains and suffers from forgetting.
To resist feature-level forgetting, a feasible solution is to freeze the pre-trained weights and append lightweight modules to encode domain-specific knowledge, e.g., prompt Jia et al. (2022). A representative work L2P Wang et al. (2022d) organizes a set of prompts as the prompt pool (i.e., $\mathbf{Pool}$) and optimizes the model via:
$\min_{\mathbf{Pool}\cup W}\sum_{({\bf x},y)\in{D^{b}}}\ell\left(W^{\top}\bar{\phi}\left({\bf x};\mathbf{Pool}\right),y\right)+\mathcal{L}_{\mathbf{Pool}}\,,$ (3)
where $\bar{\phi}\left({\bf x};\mathbf{Pool}\right)$ denotes the prompted feature representation with backbone frozen, and $\mathcal{L}_{\mathbf{Pool}}$ denotes the prompt selection loss Wang et al. (2022d). Eq. 3 dynamically retrieves instance-specific prompts to adjust the representations and learns the classifier to map the features to corresponding classes. Since the backbone weights are frozen, it can alleviate the representation drift during updating.
Discussions: Although Eq. 2 and Eq. 3 adopt different policies for model updating, both of them suffer from the feature-level forgetting. Specifically, tuning the model via Eq. 2 can encode task-specific knowledge into the embedding, but also comes with the risk of feature being fully overwritten by new domains since there is no restriction on retaining previous knowledge. Besides, although Eq. 3 freezes the pre-trained weights, the trainable prompts appended may also incur forgetting. Since the prompted feature relies on instance-specific prompt selection, optimizing the prompts for new domains still has the risk of overwriting previous ones. With the features being biased toward the latest domain, the classifier is also adjusted upon it. Consequently, the classifiers of previous tasks are incompatible with the unstable and ever-changing features, resulting in the classifier-level forgetting.
@@ 4 Duct: Dual Consolidation for Domain-Incremental Learning
Observing the risk of feature-level and classifier-level overwriting, we aim to resist forgetting in DIL from two aspects. Firstly, alleviating feature-level forgetting requires obtaining a unified embedding space at a low cost. Since sequentially updating the backbone or prompts will both result in forgetting, a proper solution is needed to make full use of all historical features and consolidate them into a unified one. Secondly, as the representation changes from task to task, the mismatch between classifier and embedding becomes more severe as the data evolves. Consequently, a classifier consolidation step is needed to calibrate them to be compatible with the embeddings.
In the following sections, we introduce the learning paradigm for representation consolidation and classifier consolidation. Lastly, we provide detailed guidelines for training and inference.
@@ 4.1 Representation Consolidation
Observing the challenge of striking a balance between all seen domains, we need to opt for another feature updating policy to avoid sequential overwriting and resist forgetting. Let us assume an ideal scenario where we can separately train each model for each incoming domain, obtaining a set of models $\{\phi_{1}(\cdot),W_{1}\},\cdots,\{\phi_{B}(\cdot),W_{B}\}$. Those models are obtained by optimizing the same PTM via Eq. 2, thus having the discriminability for each specific domain and can be seen as the ‘domain expert.’ In an oversimplified scenario where we know which domain the instance is from, we can directly utilize the corresponding expert for prediction. However, since such auxiliary information is unavailable in DIL, we need to obtain a omnipotent embedding space that suits all domains. In this way, there is no need to decide which embedding to use since the omnipotent embedding is strong enough to capture all domain-specific features.
@@
Correspondingly, we take inspiration from the model merging community Ilharco et al. (2023); Ramé et al. (2023) that an ideal model for multiple domains can be achieved by combining multiple task vectors. We denote the relative change of the embedding function as $\delta_{\phi_{i}}=\phi_{i}-\phi_{0}$, where $\phi_{0}$ is the weight of the pre-trained model. Hence, $\delta_{\phi_{i}}$ represents the relative change of the weights in the $i$-th domain, and we can build the unified embedding via:
$\phi^{m}_{i}=\phi_{0}+\alpha_{\phi}\sum_{i}\delta_{\phi_{i}}\,,$ (4)
@@
which combines the pre-trained weight with every task vector. Since all the fine-tuned embeddings start from the same pre-trained weight, Ilharco et al. (2023); Zhang & Bottou (2023) verify the task vectors share low similarity due to the semantic gap between different domains. Consequently, adding such weights enables the merged model to highlight all task-specific features, and Eq. 4 provides a feasible way to obtain the universal feature representation that suits all seen domains.
@@
Representation consolidation with task similarity: Although Eq. 4 provides a simple way to consolidate multiple task vectors to build a unified embedding, it still lacks consideration in measuring task-wise similarities. For example, if two distinct domains are more similar, highlighting those features would be more helpful for recognizing classes of corresponding domains. Hence, we introduce task-similarity ($\textbf{Sim}_{0,i}$) into the representation consolidation process and replace Eq. 4 into:
$\phi^{m}_{i}=\phi_{0}+\alpha_{\phi}\sum_{i}\textbf{Sim}_{0,i}\delta_{\phi_{i}}\,,$ (5)
@@
Specifically, since the merged embedding is built upon pre-trained weights and task vectors, we consider the similarity of pre-trained model $\phi_{0}$ and subsequent models $\phi_{i}$ to adjust the merging process. A naive way to is to directly measure the cosine similarity between $\phi_{0}$ and $\phi_{i}$. However, since pre-trained weights are with millions of dimensions, the similarity in such space is ineffective due to curse of dimensionality. To this end, we utilize the current training task as an indicator for task similarity. Specifically, for every class in $\mathcal{D}^{b}$, we utilize the backbone $\phi_{i}$ to extract class centers in its embedding space:
@@
where we utilize cosine similarity to calculate center-wise similarity $\text{sim}(\cdot,\cdot)$.
Effect of representation consolidation: Figure 1(top) visualizes the representation consolidation process, where the merged backbone can unify multiple domains. Through Eq. 5, we are able to build the joint embedding space that suits all tasks by combining them in the parameter space. Since tasks in DIL emerge one-by-one, the representation consolidation process can also be made incrementally, i.e., $\phi^{m}_{i}=\phi^{m}_{i-1}+\alpha_{\phi}\textbf{Sim}_{0,i}\delta_{\phi_{i}}$. In this way, we can get rid of the extensive memory cost and only keep at most two backbones in memory. In each training stage, we first fine-tune the model via Eq. 2 and conduct representation consolidation to aggregate the representations. In this way, we obtain a unified embedding space across all seen tasks.

### v2

@@ Abstract
Domain-Incremental Learning (DIL) involves the progressive adaptation of a model to new concepts across different domains. While recent advances in pre-trained models provide a solid foundation for DIL, learning new concepts often results in the catastrophic forgetting of pre-trained knowledge. Specifically, sequential model updates can overwrite both the representation and the classifier with knowledge from the latest domain. Thus, it is crucial to develop a representation and corresponding classifier that accommodate all seen domains throughout the learning process. To this end, we propose DUal ConsolidaTion (Duct) to unify and consolidate historical knowledge at both the representation and classifier levels. By merging the backbone of different stages, we create a representation space suitable for multiple domains incrementally. The merged representation serves as a balanced intermediary that captures task-specific features from all seen domains. Additionally, to address the mismatch between consolidated embeddings and the classifier, we introduce an extra classifier consolidation process. Leveraging class-wise semantic information, we estimate the classifier weights of old domains within the latest embedding space. By merging historical and estimated classifiers, we align them with the consolidated embedding space, facilitating incremental classification. Extensive experimental results on four benchmark datasets demonstrate Duct’s state-of-the-art performance. Code is available at https://github.com/Estrella-fugaz/CVPR25-Duct.
22footnotetext: Correspondence to: Han-Jia Ye (yehj@lamda.nju.edu.cn)
@@ 1 Introduction
Recent years have seen the rise of deep learning, demonstrating its strong potential in real-world applications [22, 16, 18, 77, 43, 44]. However, in a dynamic and ever-changing world, data often evolves in a stream format [1, 31], necessitating continuous updates with data from new domains [20]. For instance, autonomous vehicles must handle driving tasks across different seasons and weather conditions [10], and face recognition systems are supposed to recognize users despite the varying lighting conditions [82]. Correspondingly, Domain-Incremental Learning (DIL) [61] addresses this challenge by absorbing new knowledge from new data distributions while preserving existing knowledge. However, sequentially updating the model with new domains often biases the embedding [78] and classifier [81, 74] modules toward the latest domain, leading to catastrophic forgetting [41, 15] of previous knowledge. Therefore, developing methods for learning new domains without forgetting existing knowledge becomes a critical challenge for the machine learning community.
To combat catastrophic forgetting, recent works leverage strong pre-trained models (PTMs) [21] as the initialization for DIL [70, 69, 68]. PTMs’ robust feature representation abilities provide a powerful starting point, showing promising results in DIL benchmarks. These approaches freeze the pre-trained backbone to preserve existing knowledge and append lightweight modules [28, 14, 33] to capture new patterns. However, since incoming domains will also overwrite the appended modules, the representations still suffer from forgetting, leading to corresponding bias in the classifier.
In DIL, there are two primary sources of forgetting: the overwriting of features and classifiers. Continually adjusting the features to capture the latest task biases the representations toward the newest domain. Assuming the previous task involves real-world photos while the subsequent one consists of quick-draw images, the features will be overwhelmingly adjusted to highlight non-objective details. This may hinder the model from distinguishing the previous domain, resulting in forgetting. In light of this, an ideal feature space in DIL should fit all domains and equally highlight the domain-specific features for all seen tasks. However, as features become less generalizable due to overwriting, the classifier also becomes biased toward the latest domain. Since the classifier plays the pivotal role of mapping embeddings to classes, using the biased classifier can cause detrimental effects on decision-making. Consequently, it is imperative to consolidate existing knowledge at both the representation and classifier levels to resist forgetting.
@@
In this paper, we propose DUal ConsolidaTion (Duct) to address these challenges. To counter feature drift, we introduce a representation merging technique that continuously integrates historical backbones. This merged embedding consolidates task-specific information from all previous domains, providing informative features without forgetting. Consequently, we obtain non-forgettable features incrementally and resist feature-level forgetting. To map updated features to classes, we design a classifier consolidation process that merges historical classifier weights with calibrated weights. By extracting class-wise semantic relationships in the joint space, we develop a classifier transport mechanism that estimates classifiers of previous domain classes in the latest embedding space. Through consolidating features and classifiers in this coordinated way, Duct balances all seen domains and robustly resists forgetting, achieving strong performance on various benchmarks.
@@ 2 Related Work
Continual Learning/Incremental Learning aims to incorporate new knowledge within streaming data [15, 38, 67]. It can be further classified into three subcategories, i.e., task-incremental learning (TIL) [15], class-incremental learning (CIL) [38, 87], and domain-incremental learning (DIL) [61]. Among them, TIL and CIL are faced with emerging new classes and are required to learn them without forgetting. By contrast, DIL mainly focuses on the scenario where label space is fixed while incoming data involves new domains [68, 56, 42]. In continual learning, typical solutions include using data replay [8, 7, 27, 53, 51, 83, 34, 35] to review former knowledge, using knowledge distillation [23, 52, 32] or parameter regularization [29, 79, 13, 5, 6, 25] to maintain existing knowledge. Recent advances also resort to bias correction [24, 81, 2, 12, 74, 9] to rectify the model bias or expand the network structure as data evolves [55, 4, 75, 65, 85, 84, 66].
Domain-Incremental Learning with PTMs: With the rapid development of pre-training techniques, PTMs provide a strong initialization for DIL models to boost the feature generalizability [58, 69, 70, 68, 60, 49, 37, 76]. Current PTM-based DIL methods mainly resort to visual prompt tuning [28] to adjust the pre-trained features. With the pre-trained features frozen, it learns a prompt pool as external knowledge and selects instance-specific prompts to encode task information. L2P [70] utilizes the query-key matching mechanism to select instance-specific prompts, while DualPrompt [69] extends it by learning task-specific and instance-specific prompts jointly. By contrast, S-Prompts [68] learns domain-specific prompts for DIL and retrieves prompts via KNN search. To alleviate the prompt selection cost, CODA-Prompt [58] reformulates this selection step as an attention-based weighted combination process. There are also some methods that directly build classifiers upon the pre-trained features [88, 40].
Model Merging: Recent advances in pre-training have made weight interpolation among different models a useful technique in model editing [3, 19, 26, 39, 72, 73]. Most work focuses on simple averaging of multiple models, aiming to enhance model in terms of its performance on the single task and robustness to out-of-distribution data. This paper considers merging pre-trained weights and task vectors to cultivate multitask representation for all seen domains.
@@ 3.1 Domain-Incremental Learning
In domain-incremental learning, the model is faced with the data stream containing a sequence of tasks $\left\{\mathcal{D}^{1},\mathcal{D}^{2},\cdots,\mathcal{D}^{B}\right\}$, where $\mathcal{D}^{b}=\left\{\mathcal{X}_{b},\mathcal{Y}_{b}\right\}$ is the dataset of the $b$-th training task, i.e., $\mathcal{X}_{b}=\left\{{\bf x}_{i}\right\}_{i=1}^{n_{b}}$ and $\mathcal{Y}_{b}=\left\{y_{i}\right\}_{i=1}^{n_{b}}$. Each training instance ${\bf x}_{i}\in\mathbb{R}^{D}$ belongs to class $y_{i}\in Y$. This paper follows the exemplar-free setting [89, 70], i.e., during the $b$-th training stage, we can only access data in the current dataset $\mathcal{D}^{b}$. In DIL, the label space $Y$ is unchanged throughout the learning process, while the distribution of input instance keeps changing from domain to domain, i.e., $p(\mathcal{X}_{b})\neq p(\mathcal{X}_{b^{\prime}})$ for $b\neq b^{\prime}$. The target of DIL is to fit a model $f({\bf x})$ that can discriminate the classes among any seen domains:
@@
Following [69, 70, 68, 58], we assume a pre-trained Vision Transformer (ViT) [17] is available as the initialization for $f({\bf x})$. For simplicity, we decouple the network structure into the embedding function $\phi(\cdot):\mathbb{R}^{D}\rightarrow\mathbb{R}^{d}$ (i.e., the final [CLS] token) and the linear classifier $W\in\mathbb{R}^{d\times|Y|}$. In this way, the output is denoted as $f({\bf x})=W^{\top}\phi({\bf x})$, and we utilize a cosine classifier in this paper. The classifier is further decoupled into $W=[{\bf w}_{1},{\bf w}_{2},\cdots,{\bf w}_{|{Y}|}]$, where ${\bf w}_{j}$ denotes the classifier weight of the $j$-th class. Although the label space is static during the learning process, data distribution of the same class across different domains can yield significant domain gaps. Hence, we follow existing works [69, 70, 68, 58] to scale the classifier up to $W\in\mathbb{R}^{d\times b|Y|}$ when the $b$-th task emerges, and the final prediction is made by $\left(\mathop{\arg\max}_{i}{\bf w}_{i}^{\top}\phi({\bf x})\right)\textbf{ mod }|Y|$.
@@ 3.2 Baselines in Domain-Incremental Learning
In domain-incremental learning, a naive solution to handle the incoming datasets of new domains is to directly optimize the embedding and classifier:
@@
With pre-trained weights as initialization, Eq. (2) can quickly capture the discriminative features within the new domain. However, since the embedding $\phi(\cdot)$ keeps changing across different domains, it quickly loses the generalization ability to previous domains and suffers from forgetting.
To resist feature-level forgetting, a feasible solution is to freeze the pre-trained weights and append lightweight modules to encode domain-specific knowledge, e.g., prompt [28]. A representative work L2P [70] organizes a set of prompts as the prompt pool ($\mathcal{P}$) and optimizes:
$\min_{\mathcal{P}\cup W}\sum_{({\bf x},y)\in{D^{b}}}\ell\left(W^{\top}\bar{\phi}\left({\bf x};\mathcal{P}\right),y\right)+\mathcal{L}_{\mathcal{P}}\,,$ (3)
where $\bar{\phi}\left({\bf x};\mathcal{P}\right)$ denotes the prompted feature representation with backbone frozen, and $\mathcal{L}_{\mathcal{P}}$ denotes the prompt selection loss [70]. Eq. (3) dynamically retrieves instance-specific prompts to adjust the representations and learns the classifier to map the features to corresponding classes. Since the backbone weights are frozen, it can alleviate the representation drift during updating.
Discussions: Although Eq. (2) and Eq. (3) adopt different policies for model updating, both of them suffer from the feature-level forgetting. Specifically, tuning the model via Eq. (2) can encode task-specific knowledge into the embedding, but also comes with the risk of features being fully overwritten by new domains since there is no restriction on retaining previous knowledge. Besides, although Eq. (3) freezes the pre-trained weights, the trainable prompts appended may also incur forgetting. Since the prompted feature relies on instance-specific prompt selection, optimizing the prompts for new domains still has the risk of overwriting previous ones. With the features being biased toward the latest domain, the classifier is adjusted accordingly. Consequently, the classifiers of previous tasks are incompatible with the unstable and ever-changing features, resulting in the classifier-level forgetting.
@@ 4 Duct: Dual Consolidation for Domain-Incremental Learning
Observing the risk of feature-level and classifier-level overwriting, we aim to resist forgetting in DIL from two aspects. Firstly, alleviating feature-level forgetting requires obtaining a unified embedding space at a low cost. Since sequentially updating the backbone or prompts will both result in forgetting, a proper solution is needed to make full use of all historical features and consolidate them into a unified one. Secondly, as the embedding space changes from task to task, the mismatch between representation and classifier becomes more severe as the data evolves. Consequently, a classifier consolidation step is needed to calibrate them to be compatible with the embeddings.
@@ 4.1 Representation Consolidation
Observing the challenge of striking a balance between all seen domains, we need to opt for another feature updating policy to avoid sequential overwriting and resist forgetting. Let us assume an ideal scenario where we can separately train each model for each incoming domain, obtaining a set of models $\{\phi_{1}(\cdot),W_{1}\},\cdots,\{\phi_{B}(\cdot),W_{B}\}$. Those models are obtained by optimizing the same PTM via Eq. (2), thus having the discriminability for each specific domain and can be seen as the ‘domain expert.’ In an oversimplified scenario where we know which domain the instance is from, we can utilize the corresponding expert for prediction. However, since such auxiliary information is unavailable in DIL, we need to obtain an omnipotent embedding space that suits all domains. In this way, there is no need to decide which embedding to use since the omnipotent embedding is strong enough to capture all domain-specific features.
@@
Correspondingly, we take inspiration from model merging [26, 50] and cognitive-developmental theory [48] that an ideal model for multiple domains can be achieved by combining multiple task vectors. We denote the relative change of the embedding function as $\delta_{\phi_{i}}=\phi_{i}-\phi_{0}$, where $\phi_{0}$ is the weight of the pre-trained model. Hence, $\delta_{\phi_{i}}$ represents the relative change of the weights in the $i$-th domain, and we can build the unified embedding via:
$\phi^{m}_{i}=\phi_{0}+\alpha_{\phi}\sum_{k=1}^{i}\delta_{\phi_{k}}\,,$ (4)
@@
where $\alpha_{\phi},\phi^{m}_{i}$ refers to the weight of task vector $\delta_{\phi_{k}}$ and the merged backbone respectively. For task vector start from the same pre-trained weight, [26, 80] verify the task vectors share low similarity due to the semantic gap between different domains. Consequently, adding such weights enables the merged model to highlight all task-specific features, and Eq. (4) provides a feasible way to obtain the universal feature representation that suits all seen domains.
@@
Representation consolidation with task similarity: Although Eq. (4) provides a simple way to consolidate multiple task vectors to build a unified embedding, it still lacks consideration in measuring task-wise similarities. For example, if two distinct domains are more similar, highlighting those features would be more helpful for recognizing corresponding domains. Hence, we introduce task-similarity ($\textbf{Sim}_{0,i}$) into the consolidation process.
@@
To measure the similarity, a naive way is to directly measure the similarity between the embedding function of pre-trained model $\phi_{0}$ and that of the subsequent model $\phi_{i}$. However, since the weights of pre-trained model are with millions of dimensions, the similarity in such space is ineffective due to the curse of dimensionality. To this end, we utilize the current training task as an indicator of task similarity. Specifically, for every class in $\mathcal{D}^{b}$, we utilize the backbone $\phi_{i}$ to extract class centers in its embedding space:
@@
where we utilize cosine similarity to calculate center-wise similarity $\text{sim}(\cdot,\cdot)$. We then introduce task similarity into Eq. (4), and the unified embedding is denoted as:
$\phi^{m}_{i}=\phi_{0}+\alpha_{\phi}\sum_{k=1}^{i}\textbf{Sim}_{0,k}\delta_{\phi_{k}}\,,$ (7)
Effect of representation consolidation: Figure 1 (top) visualizes the representation consolidation process, where the merged backbone can unify multiple domains. Through Eq. (7), we are able to build the joint embedding space that suits all tasks by combining them in the parameter space. Since tasks in DIL emerge one-by-one, the representation consolidation process can also be made incrementally, i.e., $\phi^{m}_{i}=\phi^{m}_{i-1}+\alpha_{\phi}\textbf{Sim}_{0,i}\delta_{\phi_{i}}$. In this way, we can get rid of the extensive memory cost and only keep at most two backbones in memory. In each training stage, we first fine-tune the model via Eq. (2) and then conduct representation consolidation to aggregate the representations. In this way, we obtain a unified embedding space across all seen tasks.